OpenStudy (anonymous):

2(h-8)-h=h-16 h=8 h=-8 h=infinitely many solutions h=no solution

OpenStudy (anonymous):

please explain I have no Idea how to do this

OpenStudy (anonymous):

@GoldPhenoix

OpenStudy (anonymous):

Would you like GoldPhoenix to help you with this one or would you mind if I took a crack at it?

OpenStudy (anonymous):

Haha Please Help Yourself

OpenStudy (anonymous):

Hehe, okay. :) Please give me a few moments to scratch this out.

OpenStudy (anonymous):

first distribute the 2 through the parenthesis

OpenStudy (anonymous):

Okay Thank You Both

OpenStudy (anonymous):

And What Do You Mean Distribute the 2 @kelliegirl33

OpenStudy (anonymous):

Okay, I have arrived at the answer. Like kelliegirl133 said, you need to distribute the 2: 2(h - 8) - h and you should get 2h - 16 - h = h - 16 Does this make sense?

OpenStudy (anonymous):

Yes Thanks Guys That Makes More Sense

OpenStudy (anonymous):

now combine like terms

OpenStudy (anonymous):

So Do I Put All The H Togather

OpenStudy (anonymous):

(distributing the 2 means to multiply 2 by each term in the parenthesis. ) And yes, put the h's together.

OpenStudy (anonymous):

so would that me 4h

OpenStudy (anonymous):

you combine like terms... 2h - 16 - h = h - 16 h - 16 = h - 16 understand so far ?

OpenStudy (anonymous):

Yeah I do

OpenStudy (anonymous):

now you want to get the h's on one side and everything else on the other side

OpenStudy (anonymous):

Does this help you understand what the answer is? Do you have a guess?

OpenStudy (anonymous):

a little @tafkas77 and not really

OpenStudy (anonymous):

@kelliegirl33 For the answer (s)he is trying to get, that step isn't necessary, but I see why you did it. If (s)he did that though, (s)he'd get 0.

OpenStudy (anonymous):

Im trying to see how would i get the h on one side

OpenStudy (anonymous):

exactly....and if 0 = 0, well...you know what that means @tafkas77

OpenStudy (anonymous):

Yes, I know what the equation equaling zero means, but does (s)he? Sorry, are you a buy or a girl @Aminah14 ?

OpenStudy (anonymous):

Lol Im a Girl

OpenStudy (anonymous):

Okay, I was wondering, i have an Aunt Aminah but you can't make assumptions on sites like these. XD So do you understand what kellie is asking about an equation equalling zero?

OpenStudy (anonymous):

Lol haha Yeah I Hear You . And Not Really Lol Im Kinda Slow

OpenStudy (anonymous):

:D Hey, it's math. Sometimes this subject does that to people. Do you want to take this @kelliegirl33 or shall I?

OpenStudy (anonymous):

let me put it this way......h - 16 = h - 16 lets say h = 2 what do you get if h = 2

OpenStudy (anonymous):

14

OpenStudy (anonymous):

well 14=14

OpenStudy (anonymous):

you get -14 = -14 now what do you get if x = 3

OpenStudy (anonymous):

Close, but not quite. The equation is 2 - 16, so this answer will be negative.

OpenStudy (anonymous):

I mean if h = 3

OpenStudy (anonymous):

-13=-13

OpenStudy (anonymous):

so you see, whatever number you put in for h, the equations will equal

OpenStudy (anonymous):

Yes I do lol

OpenStudy (anonymous):

so what does that tell you

OpenStudy (anonymous):

um that if you put any number for h it will equal

OpenStudy (anonymous):

Let's look at those answer choices again. h=8 h=-8 h=infinitely many solutions h=no solution Do you see any that you can cross off the list, based on what you've seen so far?

OpenStudy (anonymous):

oh...good one tafkas

OpenStudy (anonymous):

no solution can be crossed of

OpenStudy (anonymous):

great1 (thanks kellie) :) Any others?

OpenStudy (anonymous):

I think the Answer Might Be 3

OpenStudy (anonymous):

And is that your final answer? :)

OpenStudy (anonymous):

I think your correct :)

OpenStudy (anonymous):

Beacause No Matter How Many Numbers We Try It Has Many Solutions

OpenStudy (anonymous):

PERFECT!!! :D

OpenStudy (anonymous):

good explaining tafkas

OpenStudy (anonymous):

Thanks You Guys So Much I Really Apreciate It

OpenStudy (anonymous):

thats what we are here for

OpenStudy (anonymous):

You too, kellie! You were so helpful. :) Anytime, @Aminah14 . kellie's right, that's what we do. :)

OpenStudy (anonymous):

Well Thumbs Up To U Both Good Jobs :)

OpenStudy (anonymous):

Hehe, thanks. :) You did good too.

OpenStudy (anonymous):

yes you did...good job Aminah

OpenStudy (anonymous):

Thank You :) Im Trying To Get Better Everyday Lol Ive Taken Algebra1 2 Times Already It Really sux I Guess This Is My Wake Up Call @tafkas77

OpenStudy (anonymous):

Thanks @kelliegirl33

OpenStudy (anonymous):

I think if you keep applying yourself, trying to understand it (like you did with us) you'll be just fine. Algebra can be tricky, but you rock it to your best ability as long as you keep trying to get better. I think you will if you keep showing this effort! :)

OpenStudy (anonymous):

Thanks That Really Means Alot To Me Lol

OpenStudy (anonymous):

You Guys ROCK!!!! LOL

OpenStudy (anonymous):

Ah, thanks. :) Keep it up! :)

OpenStudy (anonymous):

No Thank You Have A Nice Evening